HETAV PATEL

Grade 11,

Write all the factors of the following numbers: 36 ?????????????

Write all the factors of the following numbers: 
36 ?????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

36 = 1 × 36 36 = 2 × 18 36 = 3 × 12 36 = 4 × 9
36 = 6 × 6
Stop here, since both the factors (6) are same. Thus the factors of 36 are 1, 2, 3, 4, 6, 9, 12, 18and 36

Thanks